Grade 11Algebra

find for what value of a equation x^2 + 2x -(a^3 -3a - 3) = 0 has real roots

SuperGaming SuperheroGeek
8 Years agoGrade 11

1 Answer

venkat
8 Years ago
\Delta \geq 0 for the equation to have real roots.
\sqrt{B^2-4AC}\geq 0
here A=1 B=2 C=a2-3a-3
\sqrt{(2)^2-4(1)(a^{2}-3a-3)}\geq 0
\sqrt{4-4(a^{2}-3a-3)}\geq 0
2\sqrt{1-(a^{2}-3a-3)}\geq 0
\sqrt{-a^{2}+3a+4}\geq 0
{-a^{2}+3a+4}\geq 0
{a^{2}-3a-4}\leq 0
a^{2}-4a+a-4\leq 0
a(a-4)+1(a-4)\leq 0
(a+1)(a-4)\leq 0
-1\leq a\leq 4
\therefore a\ \epsilon \ [-1,4] \ for \ the \ given \ equation \ to \ have \ real \ roots
Please approve my answer if it helped you.